# PHANGS-JWST: Data Processing Pipeline and First Full Public Data Release

### v1

@@ Abstract
The exquisite angular resolution and sensitivity of JWST is opening a new window for our understanding of the Universe. In nearby galaxies, JWST observations are revolutionizing our understanding of the first phases of star formation and the dusty interstellar medium. Nineteen local galaxies spanning a range of properties and morphologies across the star-forming main sequence have been observed as part of the PHANGS-JWST Cycle 1 Treasury program at spatial scales of $\sim$5–50 pc. Here, we describe pjpipe, an image processing pipeline developed for the PHANGS-JWST program that wraps around and extends the official JWST pipeline. We release this pipeline to the community as it contains a number of tools generally useful for JWST NIRCam and MIRI observations. Particularly for extended sources, pjpipe products provide significant improvements over mosaics from the MAST archive in terms of removing instrumental noise in NIRCam data, background flux matching, and calibration of relative and absolute astrometry. We show that slightly smoothing F2100W MIRI data to $\farcs$9 (degrading the resolution by about 30 percent) reduces the noise by a factor of $\approx$3. We also present the first public release (DR1.0.1) of the pjpipe processed eight-band 2-21 $\mu$m imaging for all nineteen galaxies in the PHANGS-JWST Cycle 1 Treasury program. An additional 55 galaxies will soon follow from a new PHANGS-JWST Cycle 2 Treasury program.
@@ 1 Introduction
Figure 1: Color mosaics showing the JWST images for our full 19 galaxy sample. Red is a combination of the MIRI filters (F770W, F1000W, F1130W, and F2100W), green is a combination of NIRCam and MIRI (F360M and F770W), and blue is purely NIRCam (F300M and F335M). Note that in many cases, these mosaics zoom in to a region of interest within the galaxy. To see the full extent of each observation see Appendix A.
@@
Most of the PHANGS-JWST first results used data from the first four galaxies observed (NGC 628, NGC 7793, NGC 1365, and IC 5332). Observations of all 19 targets are now nearly complete, and studies are now beginning to exploit the full sample (Belfiore et al., 2023; Pathak et al., 2023). This work presents a full public data release (DR1.0.1) for the PHANGS-JWST Cycle 1 program and the associated pipeline (pjpipe  [^3] https://pjpipe.readthedocs.io/, https://doi.org/10.5281/zenodo.10458746). It also details the lessons we have learned during data acquisition for this Treasury program. The survey and much of the technical information have already been described in Lee et al. (2023), and we present here the significantly upgraded view of the reduction approach, emphasizing the (occasionally dramatic) applied changes.
@@
We have undertaken the task of building a pipeline to robustly deal with shortcomings we have discovered using the official JWST pipeline that are general to imaging observations, or in some cases more specific to observations where emission fills the field of view. These are $1/f$ noise in NIRCam data, astrometric alignment, background matching between tiles in a mosaic [^4] We will refer to these throughout as “mosaic tiles,” as opposed to the individual exposure “tiles”., and calibrating flux levels in the absence of a true background. The pipeline also addresses the reality of widely varying resolution across the JWST bands, as well as the opportunity for a small smoothing operation to dramatically improve the signal to noise at the long wavelength MIRI bands. pjpipe offers robust implementations to address these issues, and the main focus of this paper is to describe our algorithms, illustrate their application to PHANGS-JWST data, and explain how these link with the official JWST pipeline. We note that most of the issues addressed by pjpipe are not specific to PHANGS-JWST. Indeed, other programs have already made use of this pipeline to improve their data products (Peltonen et al., 2023; Chastenet, in prep).
@@
The layout of this paper is as follows. We briefly introduce the PHANGS-JWST survey in Section 2, before detailing our data processing steps in Section 3, noting differences between this and our early science efforts (Lee et al., 2023). We provide details of our quality assurance checks in Section 4, highlighting remaining known issues in Section 4.3. We compare to our early science mosaics and those obtained from MAST in Section 5. A description of the output files for the data release is given in Section 6. We summarise in Section 7, and present an atlas of our 19 galaxies in Appendix A. Our data (both combined mosaics and individual tiles for PSF photometry) are available publicly at https://archive.stsci.edu/hlsp/phangs-jwst.
@@ 2 Data Overview
PHANGS-JWST is a JWST Cycle 1 Treasury Program, building upon existing PHANGS efforts to map nearby ($D$ < 20 Mpc) galaxies with Atacama Large Millimeter/submillimeter Array (ALMA; Leroy et al., 2021) CO observations, Multi Unit Spectroscopic Explorer (MUSE; Emsellem et al., 2022) optical integral field spectroscopy, and high-resolution Hubble Space Telescope (HST; Lee et al., 2022) optical imaging. Specifically, this Cycle 1 Treasury (program 2107, PI: Lee) has observed the 19 main-sequence galaxies that form the PHANGS-MUSE subset of the full sample, which have been observed with JWST between 2022–06–06 and 2023–07–11555An up-to-date observing log can be found at https://www.stsci.edu/cgi-bin/get-visit-status?id=2107&markupFormat=html&observatory=JWST&pi=1.
@@ 3 Data Processing
In Lee et al. (2023), we presented our processing methodology for the four galaxies observed during the first month and a half of our program. Since then, we have seen significant improvements in a number of directions. Firstly, the reference files (for this release, we use reference files jwst_1147.pmap [^6] https://jwst-crds.stsci.edu/) have improved as the in-flight performance of JWST and its instruments have been better understood. Secondly, improvements in the official JWST pipeline [^7] https://github.com/spacetelescope/jwst (Bushouse et al., 2023) have fixed a number of bugs and improved many of the underlying reduction steps (for our DR1.0.1 reprocessing, we use pipeline version 1.12.4). Finally, we have identified and understood several issues associated with observations of extended, bright sources that fill the field of view and developed robust strategies for many of these. This paper will focus on this last point, but we note our latest data processing represents a large improvement both over our original data release (e.g., improved flux calibration, background matching and astrometry), as well as the mosaics that are currently available in the MAST archive (see Sect. 5).
Our pipeline (pjpipe) wraps around the official JWST pipeline (Bushouse et al., 2023), adding a number of bespoke steps not present in the official pipeline, which we show schematically in Figure 2. We make pjpipe publicly available at https://github.com/phangsTeam/pjpipe, which also includes the configuration required to reproduce our reduction. We show the main output of pjpipe, the mosaics, in Appendix A. For users who wish to use pjpipe for their science, documentation and examples of configuration files are available at https://pjpipe.readthedocs.io/en/. As an open-source project, we also encourage users to contribute bugfixes and new features. By making our code fully available, we hope that some of the general techniques we have developed for our image processing will also feed back into improvements in the official JWST pipeline [^8] Many of the pjpipe algorithms have been presented at the “Improving JWST Data Workshop,” and the materials and discussions from that workshop are available at https://outerspace.stsci.edu/display/JEA/Improving+JWST+Data+Products+Workshop.. A current list of known issues with the JWST pipeline is maintained at https://jwst-docs.stsci.edu/jwst-calibration-pipeline-caveats/known-issues-with-jwst-data-products (some of which are dealt with by pjpipe).
@@ 3.1 Level One Processing
The Level One stage of the pipeline transforms uncalibrated images to count rate images via a number of steps (e.g., flagging bad pixels, subtracting reference pixels, and fitting slopes to ramps). The details for each step are given at https://jwst-pipeline.readthedocs.io/en/latest/jwst/pipeline/calwebb_detector1.html [^9] For the specific documentation for this version, see https://doi.org/10.5281/zenodo.8436689. We essentially keep this step at the observatory-recommended defaults, although we process individual exposures (tiles) from each dither group in observation sequence rather than simultaneously. Processing the observations sequentially allows us to use the output files from the previous observation to identify persistence, i.e., faint residual images caused by previous exposure to bright sources appearing in subsequent integrations. Then data affected by persistence can be flagged in the subsequent persistence step. We also attempt to mitigate saturation by setting suppress_one_group to False in the ramp fitting step so that a flux can be obtained for any pixels where the first group is not saturated.
@@ 3.2 Level Two Processing
In this stage, we combine into a single master background frame the MIRI background observations in image space. A background estimate is calculated per-pixel after applying sigma-clipping and combining all dithers executed during the off observation. In the case of PHANGS-JWST, the MIRI background observations are obtained in parallel during the NIRCam science observations and as a result, the dither used is also the one executed to obtain the NIRCam images. When constructing the background image, we use a relatively strict sigma-clip of $1.5\sigma$. Our background fields occasionally contain clear foreground stars or background galaxies and this strict tolerance limits the number of negative artifacts seen in our calibrated images, although in a few cases unavoidable artifacts persist. For our larger mosaics, we expect the overall noise level to be slightly lower because we acquire background observations for every NIRCam position, meaning we generate a background from more observations.
@@
When constructing MIRI background images, proximity of the background to the on-source image in time appears to be a critical factor. For PHANGS-JWST, the initial NIRCam observations, and hence the MIRI background images, failed for IC 5332, and the successful NIRCam science and MIRI background observations were only obtained months later. Another PHANGS-JWST target, NGC 7496 was observed immediately prior to IC 5332. Figure 3 compares the results subtracting backgrounds obtained close in position but in not time (i.e., using the later IC 5332 for the background) and close in time but not in position (i.e., using NGC 7496). This shows that using the MIRI backgrounds obtained months after the MIRI science observations led to large-scale artifacts in the final mosaics, highlighting the temporal variability of the MIRI backgrounds. In the current release, we use the NGC 7496 “off” observations that ended 14 minutes before the IC 5332 science observations started to construct the background for IC 5332. For pointings where the background observations fail, we would therefore recommend re-observing both science and background given the clear time variability of the MIRI background.
@@ 3.5 Background Matching
[…]
@@
We therefore take a different approach. For each pixel in the overlap region between a pair of images, including those that contain signal from the source, we calculate the difference between values at that pixel in the different images. Then we take the median of the pixel-by-pixel differences to calculate the average offset between these two images. To do this, we first reproject all images to a common pixel grid and then calculate the median difference of all overlapping pixels. When calculating these median differences between pairs of tiles, we use sigma-clipping to reject outlying differences. We then additively adjust the background level of the images being considered to minimize the differences among all image pairs being considered in a least-squares sense, weighting each image pair by the root mean square (RMS) scatter of the difference histograms for each tile overlap. To avoid including noisy and untrustworthy overlaps between images, we do not include any image pairs with a total pixel overlap less than 0.2 percent the average number of pixels in an overlap region. Nor do we include image pairs that do not overlap by at least 25 percent along at least one axis of the image or image pairs that have RMS values more than $2\sigma$ greater than the average RMS value of all the other image pair differences. These cuts typically remove the small diagonal overlaps that occur in the centers of the $2\times 2$ mosaics or similar (or the centers of the four NIRCam short chips). These choices mirror those used by Montage (Jacob et al., 2010) for their background matching step.
@@ 3.6 NIRCam Destriping
Note that for our initial reduction of the first four galaxies (presented in Lee et al., 2023, and associated papers), we used a robust principal component analysis (PCA) method. Further testing has revealed that this method is extremely sensitive to source masking, and using a median filter yielded more robust results. We do believe that PCA may ultimately represent the optimal strategy when there is limited extended emission in a tile, but our data contain too much emission for the PCA to work successfully.

### v3

@@ Abstract
The exquisite angular resolution and sensitivity of JWST is opening a new window for our understanding of the Universe. In nearby galaxies, JWST observations are revolutionizing our understanding of the first phases of star formation and the dusty interstellar medium. Nineteen local galaxies spanning a range of properties and morphologies across the star-forming main sequence have been observed as part of the PHANGS-JWST Cycle 1 Treasury program at spatial scales of $\sim$5–50 pc. Here, we describe pjpipe, an image processing pipeline developed for the PHANGS-JWST program that wraps around and extends the official JWST pipeline. We release this pipeline to the community as it contains a number of tools generally useful for JWST NIRCam and MIRI observations. Particularly for extended sources, pjpipe products provide significant improvements over mosaics from the MAST archive in terms of removing instrumental noise in NIRCam data, background flux matching, and calibration of relative and absolute astrometry. We show that slightly smoothing F2100W MIRI data to $\farcs$9 (degrading the resolution by about 30 percent) reduces the noise by a factor of $\approx$3. We also present the first public release (DR1.1.0) of the pjpipe processed eight-band 2-21 $\mu$m imaging for all nineteen galaxies in the PHANGS-JWST Cycle 1 Treasury program. An additional 55 galaxies will soon follow from a new PHANGS-JWST Cycle 2 Treasury program.
@@
facilities: JWST.software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), jwst (Bushouse et al., 2023), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), photutils (Bradley et al., 2022), SAOImageDS9 (Joye & Mandel, 2003), scipy (Virtanen et al., 2020).
@@ 1 Introduction
Figure 1: Color mosaics showing the JWST images for our full 19 galaxy sample developed to support a 2024 January press release, https://webbtelescope.org/contents/news-releases/2024/news-2024-105. Red is a combination of the MIRI filters (F770W, F1000W, F1130W, and F2100W), green is a combination of NIRCam and MIRI (F360M and F770W), and blue is purely NIRCam (F300M and F335M). F200W imaging has also been obtained, but is not included in the color composites. Note that in many cases, these mosaics are zoomed, and so do not show the full extent of coverage of the galaxy. For the full coverage of each observation, see Appendix A.
@@
Most of the PHANGS-JWST first results used data from the first four galaxies observed (NGC 628, NGC 7793, NGC 1365, and IC 5332). Observations of all 19 targets are now complete, and studies are beginning to exploit the full sample (Belfiore et al., 2023; Pathak et al., 2023). This work presents a full public data release (DR1.1.0333Some issues were found with the astrometry in NIRCam observations of NGC 3627 and have been updated to a DR1.1.1. Other data remains unchanged.) for the PHANGS-JWST Cycle 1 program and the associated pipeline (pjpipe  [^4] https://pjpipe.readthedocs.io/, https://doi.org/10.5281/zenodo.10458746). It also details the lessons we have learned during data acquisition for this Treasury program. The survey and much of the technical information have already been described in Lee et al. (2023), and we present here the significantly upgraded view of the reduction approach, emphasizing the (occasionally dramatic) applied changes.
@@
We have undertaken the task of building a pipeline to robustly deal with shortcomings we have discovered using the official JWST pipeline that are general to imaging observations, or in some cases more specific to observations where emission fills the field of view. These are $1/f$ noise in NIRCam data, astrometric alignment, background matching between tiles in a mosaic [^5] We will refer to these throughout as “mosaic tiles,” as opposed to the individual single exposure “tiles” (in the case of the NIRCam short images, these are the individual detector images, rather than the combination of the multiple detectors). These mosaic tiles are formed by the combination of all (four) of the dithers in a single pointing (for the NIRCam short images, this combines the four individual detectors as well, for a total of 16 input exposures to a mosaic tile)., and calibrating flux levels in the absence of a true background. The pipeline also addresses the reality of widely varying resolution across the JWST bands, as well as the opportunity for a small smoothing operation to dramatically improve the signal to noise at the long wavelength MIRI bands. pjpipe offers robust implementations to address these issues, and the main focus of this paper is to describe our algorithms, illustrate their application to PHANGS-JWST data, and explain how these link with the official JWST pipeline. We note that most of the issues addressed by pjpipe are not specific to PHANGS-JWST. Indeed, other programs have already made use of this pipeline to improve their data products (Peltonen et al., 2023; Chastenet, in prep).
@@
The layout of this paper is as follows. We briefly introduce the PHANGS-JWST survey in Section 2, before detailing our data processing steps in Section 3, noting differences between this and our early science efforts (Lee et al., 2023). We provide details of our quality assurance checks in Section 4, highlighting remaining known issues in Section 4.3. We compare to our early science mosaics and those obtained from MAST in Section 5. A description of the output files for the data release is given in Section 6. We summarise in Section 7, and present an atlas of our 19 galaxies in Appendix A. Our data (both combined mosaics and individual tiles for PSF photometry) are available publicly at https://archive.stsci.edu/hlsp/phangs/phangs-jwst.
@@ 2 Data Overview
PHANGS-JWST is a JWST Cycle 1 Treasury Program, building upon existing PHANGS efforts to map nearby ($D$ < 20 Mpc) galaxies with Atacama Large Millimeter/submillimeter Array (ALMA; Leroy et al., 2021) CO observations, Multi Unit Spectroscopic Explorer (MUSE; Emsellem et al., 2022) optical integral field spectroscopy, and high-resolution Hubble Space Telescope (HST; Lee et al., 2022) optical imaging. Specifically, this Cycle 1 Treasury (program 2107, PI: Lee) has observed the 19 main-sequence galaxies that form the PHANGS-MUSE subset of the full sample, which have been observed with JWST between 2022–06–06 and 2024–02–06666An up-to-date observing log can be found at https://www.stsci.edu/cgi-bin/get-visit-status?id=2107&markupFormat=html&observatory=JWST&pi=1.
@@
s $\mu$Jy MJy sr-1
NIRCam
F200W 1202.5 16 0.039 0.071
F300M 386.5 4 0.084 0.052
F335M 386.5 4 0.092 0.050
F360M 429.5 8 0.107 0.054
MIRI
F770W 88.8 4 1.028 0.13
F1000W 122.1 4 1.253 0.13
F1130W 310.8 4 2.637 0.24
F2100W 321.9 4 5.152 0.27
Table 1: Overview of the observational setup for our JWST filters. We use a four-point dither pattern in all cases, so the average tile overlap is an indication of how many repeats we perform in the same area. Exposure times and overlaps are from Lee et al. (2023), but sensitivities have been updated using the same method presented in Lee et al. (2023), reflecting the new data processing.
For full details of the survey, we refer the reader to Lee et al. (2023) and we briefly summarize details relevant to this paper here. The observations cover eight bands, spanning 2–21 µm, using both NIRCam (F200W, F300M, F335M, F360M) and MIRI (F770W, F1000W, F1130W, F2100W). Our observations are primarily mosaics, with a maximum 2$\times$2 tiling pattern. Because our targets are large on the sky, even these mosaics are often not sufficient to cover the full galaxy, though the coverage always captures most of the infrared emission and star formation activity from the galaxy (Figure 1), and we place the JWST footprint to overlap existing data. Details of the observations are given in Table 1. We use a minimum of four dithers per integration to fully sample the PSF, up to a maximum of sixteen overlapping dithers in the case of F200W. To maximize the observing efficiency, we observed off positions for the MIRI data in parallel with our NIRCam science observations. A collage of our mosaics is shown in Figure 1. A rich tapestry of structure is clearly visible, from extended dust emission to point-like, embedded stellar clusters.
@@ 3 Data Processing
In Lee et al. (2023), we presented our processing methodology for the four galaxies observed during the first month and a half of our program. Since then, we have seen significant improvements in a number of directions. Firstly, the reference files (for this release, we use reference files jwst_1201.pmap [^7] https://jwst-crds.stsci.edu/) have improved as the in-flight performance of JWST and its instruments have been better understood. Secondly, improvements in the official JWST pipeline [^8] https://github.com/spacetelescope/jwst (Bushouse et al., 2023) have fixed a number of bugs and improved many of the underlying reduction steps (for our DR1.1.0 reprocessing, we use pipeline version 1.13.4). Finally, we have identified and understood several issues associated with observations of extended, bright sources that fill the field of view and developed robust strategies for many of these. This paper will focus on this last point, but we note our latest data processing represents a large improvement both over our original data release (e.g., improved flux calibration, background matching and astrometry), as well as the mosaics that are currently available in the MAST archive (see Sect. 5).
Our pipeline (pjpipe) wraps around the official JWST pipeline (Bushouse et al., 2023), adding a number of bespoke steps not present in the official pipeline, which we show schematically in Figure 2. We make pjpipe publicly available at https://github.com/phangsTeam/pjpipe, which also includes the configuration required to reproduce our reduction. We show the main output of pjpipe, the mosaics, in Appendix A. For users who wish to use pjpipe for their science, documentation and examples of configuration files are available at https://pjpipe.readthedocs.io/en/. As an open-source project, we also encourage users to contribute bugfixes and new features. By making our code fully available, we hope that some of the general techniques we have developed for our image processing will also feed back into improvements in the official JWST pipeline [^9] Many of the pjpipe algorithms have been presented at the “Improving JWST Data Workshop,” and the materials and discussions from that workshop are available at https://outerspace.stsci.edu/display/JEA/Improving+JWST+Data+Products+Workshop.. A current list of known issues with the JWST pipeline is maintained at https://jwst-docs.stsci.edu/known-issues-with-jwst-data (some of which are dealt with by pjpipe).
@@ 3.1 Level One Processing
The Level One stage of the pipeline transforms uncalibrated images to count rate images via a number of steps (e.g., flagging bad pixels, subtracting reference pixels, and fitting slopes to ramps). The details for each step are given at https://jwst-pipeline.readthedocs.io/en/latest/jwst/pipeline/calwebb_detector1.html [^10] For the specific documentation for this version, see https://doi.org/10.5281/zenodo.8436689. We essentially keep this step at the observatory-recommended defaults, although we process individual exposures (tiles) from each dither group in observation sequence rather than simultaneously. Processing the observations sequentially allows us to use the output files from the previous observation to identify persistence, i.e., faint residual images caused by previous exposure to bright sources appearing in subsequent integrations. Then data affected by persistence can be flagged in the subsequent persistence step. In our observations, we have not noticed persistence occurring from previous observations, but this could be a potential issue for observations immediately following, for example, observations of solar system planets. We also attempt to mitigate saturation by setting suppress_one_group to False in the ramp fitting step so that a flux can be obtained for any pixels where the first group is not saturated.
@@ 3.2 Level Two Processing
In this stage, we combine into a single master background frame the MIRI background observations in image space. A background estimate is calculated per-pixel after applying sigma-clipping and combining all dithers executed during the off observation. In the case of PHANGS-JWST, the MIRI background observations are obtained in parallel during the NIRCam science observations and as a result, the dither used is also the one executed to obtain the NIRCam images. When constructing the background image, we use a relatively strict sigma-clip of $1.5\sigma$. Our background fields occasionally contain clear foreground stars or background galaxies and this strict tolerance limits the number of negative artifacts seen in our calibrated images, although in a few cases unavoidable artifacts persist. These are extended background galaxies that the dither pattern fails to remove, and appear as negative spots in our background images. Work is ongoing to mitigate for this in this step (e.g. by masking sources and interpolating over missing data), but our current data occasionally contain these artifacts. For our larger mosaics, we expect the overall noise level to be slightly lower because we acquire background observations for every NIRCam position, meaning we generate a background from more observations.
@@
When constructing MIRI background images, proximity of the background to the on-source image in time appears to be a critical factor. For PHANGS-JWST, the initial NIRCam observations, and hence the MIRI background images, failed for IC 5332, and the successful NIRCam science and MIRI background observations were only obtained months later. Another PHANGS-JWST target, NGC 7496 was observed immediately prior to IC 5332. Figure 3 compares the results subtracting backgrounds obtained close in position but not in time (i.e., using the later IC 5332 for the background) and close in time but not in position (i.e., using NGC 7496). We perform a sigma-clipped RMS measurement on these three images, using a source mask created from the “close in time, not in position” backgrounds. We obtain RMS=[0.80, 0.46, 0.36] MJy sr-1 in these three cases. This shows that using the MIRI backgrounds obtained months after the MIRI science observations led to large-scale artifacts in the final mosaics, highlighting the temporal variability of the MIRI backgrounds. In the current release, we use the NGC 7496 “off” observations that ended 14 minutes before the IC 5332 science observations started to construct the background for IC 5332. For pointings where the background observations fail, we would therefore recommend re-observing both science and background given the clear time variability of the MIRI background.
@@ 3.5 Background Matching
Particularly for MIRI data, the background levels can vary substantially, even after subtracting the background calculated in the Level Two processing. This is mainly driven by instrumental backgrounds (primarily the telescope thermal background at longer wavelengths, and scattered light at shorter wavelengths), which we observe to vary at the $\sim$0.1 MJy sr-1 level between dithers even within a single mosaic tile. These must be corrected for in order to provide consistent level matching between individual observations, and tiles in a mosaic.
@@
We therefore take a different approach. For each pixel in the overlap region between a pair of images, including those that contain signal from the source, we calculate the difference between values at that pixel in the different images. Then we take the median of the pixel-by-pixel differences to calculate the average offset between these two images. To do this, we first reproject all images to a common pixel grid and then calculate the median difference of all overlapping pixels. When calculating these median differences between pairs of tiles, we use sigma-clipping to reject outlying differences. We then adjust the background level of the images being considered, in order to minimize the differences among all image pairs being considered. This is done using least-squares minimization, weighting each image pair by the root mean square (RMS) spread of the difference histograms for each tile overlap. To avoid including noisy and untrustworthy overlaps between images, we do not include any image pairs with a total pixel overlap less than 0.2 percent the average number of pixels in an overlap region. We also do not include image pairs that do not overlap by at least 25 percent along at least one axis of the image, or image pairs that have RMS values that are more than $2\sigma$ greater than the average RMS value of all the other image pair differences (i.e. very noisy difference histograms, which often occur when there are few overlapping pixels). These cuts typically remove the small diagonal overlaps that occur in the centers of the $2\times 2$ mosaics or similar (or the centers of the four NIRCam short chips). These choices mirror those used by Montage (Jacob et al., 2010) for their background matching step.
@@ 3.6 NIRCam Destriping
Note that for our initial reduction of the first four galaxies (presented in Lee et al., 2023, and associated papers), we used a robust principal component analysis (PCA) method. Further testing has revealed that this method is extremely sensitive to source masking, and using a median filter yielded more robust results. Particularly, our testing with the PCA method often showed imprints of the source in the noise model, indicating that we are removing real emission. The same is not true in the case of the median filter. We do believe that PCA may ultimately represent the optimal strategy when there is limited extended emission in a tile, but our data contain too much emission for the PCA to work successfully.